Question #1: Solving for the Goods Market Suppose an economy can be described by the following set of equations: C = 160 +(0.(g) What is the marginal propensity to consume (mpc)? (h) What is the marginal propensity to save (mps)? (i) Calculate the muuestion #1:Solving for the Goods MarketSuppose an economy can be described by the following set of equations:C = 160+(0.6)YDI = 150G = 150T = 100YD= Y -TNX = 0 (g) What is the marginal propensity to consume (mpc)?(h) What is the marginal propensity to save (mps)?(i) Calculate the multiplier. [Hint: The multiplier = 1/(1-mpc)](j) Calculate the level of autonomous spending. [Hint: Autonomous spending = C0+ I + G –(mpc x T)]
Answer #1

= 160 to.blo (9) Marginal propensity to consume (MPC) I MPC = 6.6 (6) marginal propensity to save (ups). MPS = 1 - upc = 1- 0